# CLASP: An online learning algorithm for Convex Losses And Squared Penalties

### v1

@@ 3 The CLASP algorithm
Here, $\nabla f_{t}(x_{t})$ represents a sub-gradient of $f_{t}$ at $x_{t}$. The projection step in CLASP is onto the set ${\mathcal{K}}_{t}={\mathcal{K}}\cap C_{t}$, with $C_{t}=\left\{x\in\mathbb{R}^{n}\,\colon\,g_{t}(x)\leq 0\right\}$ (as defined in (2)). Note that the set ${\mathcal{K}}_{t}$ is non-empty, as per Assumption 3, and that this projection, typically realized as a convex optimization problem, sets CLASP apart from the projection-free methods in Section 1.1. The step-sizes $\eta_{t}$ are chosen according to whether the loss functions are convex or strongly convex and are specified further ahead in Section 5. We write $\widehat{x}_{t+1}=x_{t}-\eta_{t}\nabla f_{t}(x_{t})$, and rewrite CLASP as $x_{t+1}=p_{{\mathcal{K}}_{t}}(\widehat{x}_{t+1})$ for $t\geq 1$.
@@ A.4 Additional Numerical Experiments
In this section, we present additional numerical experiments. In particular, a version of the Online Linear Regression experiment with additional rounds, and an experiment for Online Support Vector Machine. The results were obtained by averaging over $100$ trials and reported with a $95\%$ confidence interval.
@@ A.4.2 Online Support Vector Machine
or, in terms of the COCO framework, the revealed loss function is $f_{t}(x)=\frac{1}{2}\|w\|^{2}$ and the revealed constraint function is $g_{t}(x)=-v_{t}\left(w^{T}u_{t}-b\right)+1$. For this experiment, we use the real-world dataset about wine quality from their physicochemical properties (for the details about the dataset, see (Cortez et al., 2009)) [^1] This dataset is released under a Creative Commons Attribution 4.0 International (CC BY 4.0) license.. The dataset contains 6497 samples, each sample contains $P=11$ features, and the quality of the wine is classified between 1 and 9. Some of the features were in $g/dm^{3}$, while others were in $mg/dm^{3}$. We ensured that all density features were expressed in $g/dm^{3}$. In this experiment, we consider the binary classification setting, where we label with $1$ the wine samples with quality equal to or greater than 7, and label with 1-1 the remaining. Thus, we want our classifier to distinguish between high-quality and low-quality wines. In each trial, we reshuffle the dataset so that the order of samples in each trial is always different. The vector $x_{1}\in\mathbb{R}^{P+1}$ for all algorithms was initialized such that $x_{i}\sim U(-1,1)$, for $1\leq i\leq P+1$. The results were obtained by averaging over $100$ trials and reported with a $95\%$ confidence interval.

### v2

@@ 3 The CLASP algorithm
Here, $\nabla f_{t}(x_{t})$ represents a sub-gradient of $f_{t}$ at $x_{t}$. The projection step in CLASP is onto the set ${\mathcal{K}}_{t}={\mathcal{K}}\cap C_{t}$, with $C_{t}=\left\{x\in\mathbb{R}^{n}\,\colon\,g_{t}(x)\leq 0\right\}$ (as defined in (2)). Note that the set ${\mathcal{K}}_{t}$ is non-empty, as per Assumption 3, and that this projection, typically realized as a convex optimization problem, sets CLASP apart from the projection-free methods in Section $1.$ The step-sizes $\eta_{t}$ are chosen according to whether the loss functions are convex or strongly convex and are specified further ahead in Section 5. We write $\widehat{x}_{t+1}=x_{t}-\eta_{t}\nabla f_{t}(x_{t})$, and rewrite CLASP as $x_{t+1}=p_{{\mathcal{K}}_{t}}(\widehat{x}_{t+1})$ for $t\geq 1$.
@@ A.4 Additional Numerical Experiments
In this section, we present additional numerical experiments. In particular, a version of the Online Linear Regression experiment with additional rounds, and an experiment for Online Support Vector Machine. The results were obtained by averaging over 100100 trials and reported with a $95\%$ confidence interval.
@@ A.4.2 Online Support Vector Machine
or, in terms of the COCO framework, the revealed loss function is $f_{t}(x)=\frac{1}{2}\|w\|^{2}$ and the revealed constraint function is $g_{t}(x)=-v_{t}\left(w^{T}u_{t}-b\right)+1$. For this experiment, we use the real-world dataset about wine quality from their physicochemical properties (for the details about the dataset, see (Cortez et al., 2009)) [^1] This dataset is released under a Creative Commons Attribution 4.0 International ($C$ BY 4.0) license.. The dataset contains 6497 samples, each sample contains $P=11$ features, and the quality of the wine is classified between 1 and 9. Some of the features were in $g/dm^{3}$, while others were in $mg/dm^{3}$. We ensured that all density features were expressed in $g/dm^{3}$. In this experiment, we consider the binary classification setting, where we label with 11 the wine samples with quality equal to or greater than 7, and label with 1-1 the remaining. Thus, we want our classifier to distinguish between high-quality and low-quality wines. In each trial, we reshuffle the dataset so that the order of samples in each trial is always different. The vector $x_{1}\in\mathbb{R}^{P+1}$ for all algorithms was initialized such that $x_{i}\sim U(-1,1)$, for $1\leq i\leq P+1$. The results were obtained by averaging over 100100 trials and reported with a $95\%$ confidence interval.